A sky diver jumps from an airplane at 5000 meters and falls for one minute at an average velocity of 51 m/sec. How much does he have to fall after his parachute opens?
1
Expert's answer
2017-11-30T15:55:07-0500
He has to fall following distance:
s = 5000 -51* 60 = 5000 - 3060 = 1940 m